if 8ƒ(x)=6ƒ(1/x)=x+5and y=x^2ƒ(x)then find the value of of dy/dx at x=1
plz sir make me understand this question step by step
note x^2 means x square

Grade:

1 Answers

Ramesh V
14 years ago

I'll take the question to be :

                                                     8ƒ(x)+6ƒ(1/x)=x+5    ........(1)

now put 1/x for x , we get ->    8ƒ(1/x)+6ƒ(x)=(1/x)+5   ....(2)

on solving, we'll get f(x) = { 4x -(3/x)+5 } / 14

so Y = x2/14* (4x -(3/x)+5 )

dy/dx = (12x2 + 10x -3)/14

dy/dx at x=1 is 19/14